OpenStudy (anonymous):

What's the equation? y is 20 less than the product of five and half of x.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

what have you tried to do?

OpenStudy (anonymous):

um y-20 + 5 1/2x?

OpenStudy (solomonzelman):

y is 20 less than the product of five and half of x. y is 20 less than 5 \(\large\color{slate}{ \cdot }\) \(\large\color{slate}{ \frac{1}{2}x }\) ( y is 20 less than this product, -- that means that if you add 20 to the y, then it will be equivalent to this product. ) y + 20 = 5 \(\large\color{slate}{ \cdot }\) \(\large\color{slate}{ \frac{1}{2}x }\)

OpenStudy (solomonzelman):

or, you can say, y = 5 \(\large\color{slate}{ \cdot }\) \(\large\color{slate}{ \frac{1}{2}x }\) - 20

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

yw